# Causal Clustering for Conditional Average Treatment Effects Estimation and Subgroup Discovery

### v1

@@ IV-D Adversarial Simulation Design
For this simulation design, we evaluate our framework as a plug-in ad-hoc CATE estimator on an adversarial structure with no true underlying distinct cluster subpopulations. We follow the simulation frameworks proposed by (athey2019generalized, ), (kunzel2019metalearners, ), and (friedberg2020local, ) for evaluating our framework. We generate $n$ samples with $p$ features by drawing $X\sim U([0,1]^{p})$. We fix the treatment assignment propensity as $e(X)=0.5$ for all $X$, and set the baseline effect of the response to zero across features, i.e., $\forall i:Y_{i}^{(0)}=0$. To assess robustness, we vary the sample size for training data in increments of 200, so that $n\in\{800,1000,1200\}$, while maintaining a fixed feature dimension $p=20$. Additionally, we introduce iid Gaussian noise $\varepsilon_{i}\sim\mathcal{N}(0,\sigma)$ to each observed outcome $Y_{i}^{obs}$, with the standard deviation of noise $\sigma\in\{1,2,3,4\}$. Model evaluation is conducted on a consistent unseen set of 2000 test samples (where only features are observed). We evaluate the excess risk of our clustering methodology against the underlying learner using the PEHE and Excess Risk metric:
@@
Following (athey2019generalized, ) and (friedberg2020local, ), we adopt this challenging response function due to its smooth, highly non-linear nature and the presence of numerous noisy irrelevant covariates and exogenous noise, in addition to the fact that no finite number of hard clusters can fully capture the true CATE without loss i.e., hard discrete assignment constraints handicap us versus continuous/soft assignments. A visualization of its non-trivial smooth structure is presented in Figure 3. We present our simulation results for the excess risk incurred by the clustering step in Figure 2.
@@ IV-E Observational Dataset Analysis: Synthea Emergency Health Records of Viral Sinusitis Treatment
For this section, we conduct causal clustering analysis based off the Synthea dataset (walonoski2018synthea, ), of patient emergency health records (EHR) and outcomes based off their real life counterparts.
@@
Following the setup and pre-processing of (lee2025accurate, ), the dataset includes timestamps of visits for a curated list of 7 other pre-treatment comorbidities: asthma, allergy, anemia, flu, hypertension, obesity, pregnancy to serves as irrelevant noise features. An additional continuous feature is derived for each patient: the average minimum distance between the two conditions. We then ran our framework and chose a cluster size of $k=3$ via the eigengap method. We present the violin plots of features per cluster in Figure 5.

### v2

@@ IV-D Adversarial Simulation Design
For this simulation design, we evaluate our framework as a plug-in ad-hoc CATE estimator on an adversarial structure with no true underlying distinct cluster subpopulations. We follow the simulation frameworks proposed by [5], [9], and [6] for evaluating our framework. We generate $n$ samples with $p$ features by drawing $X\sim U([0,1]^{p})$. We fix the treatment assignment propensity as $e(X)=0.5$ for all $X$, and set the baseline effect of the response to zero across features, i.e., $\forall i:Y_{i}^{(0)}=0$. To assess robustness, we vary the sample size for training data in increments of 200, so that $n\in\{800,1000,1200\}$, while maintaining a fixed feature dimension $p=20$. Additionally, we introduce iid Gaussian noise $\varepsilon_{i}\sim\mathcal{N}(0,\sigma)$ to each observed outcome $Y_{i}^{obs}$, with the standard deviation of noise $\sigma\in\{1,2,3,4\}$. Model evaluation is conducted on a consistent unseen set of 2000 test samples (where only features are observed). We evaluate the excess risk of our clustering methodology against the underlying learner using the PEHE and Excess Risk metric:
@@
Following [5] and [6], we adopt this challenging response function due to its smooth, highly non-linear nature and the presence of numerous noisy irrelevant covariates and exogenous noise, in addition to the fact that no finite number of hard clusters can fully capture the true CATE without loss i.e., hard discrete assignment constraints handicap us versus continuous/soft assignments. A visualization of its non-trivial smooth structure is presented in Figure 3. We present our simulation results for the excess risk incurred by the clustering step in Figure 2.
@@ IV-E Observational Dataset Analysis: Synthea Emergency Health Records of Viral Sinusitis Treatment
For this section, we conduct causal clustering analysis based off the Synthea dataset [37], of patient emergency health records (EHR) and outcomes based off their real life counterparts.
@@
Following the setup and pre-processing of [38], the dataset includes timestamps of visits for a curated list of 7 other pre-treatment comorbidities: asthma, allergy, anemia, flu, hypertension, obesity, pregnancy to serves as irrelevant noise features. An additional continuous feature is derived for each patient: the average minimum distance between the two conditions. We then ran our framework and chose a cluster size of $k=3$ via the eigengap method. We present the violin plots of features per cluster in Figure 5.